A halfback is initially running south at a speed of 11.2 m/s, quickly cuts to his right, and 0.545 s later is running 60.0° west of south at 11.2 m/s. Find his average acceleration.

Magnitude:

direction ________ north of west.
Answer #1

Here,

let east is the positive x direction and north is the positive y direction

for the initial velocity

u = -11.2 j m/s

final velocity

v = 11.2 * (- sin(60) i - j * cos(60))

Now, for the average acceleration

a = (v - u)/time

a = ( 11.2 * (- sin(60) i - j * cos(60))- (-11.2j))/0.545

a = -17.8 i + 10.3 j

magnitude of acceleration = sqrt(17.8^2 + 10.3^2) = 20.6 m/s^2

the magnitude of acceleration is 20.6 m/s^2

direction = atan(10.3/17.8) degree north of west

the direction = 30.1 degree north of west

the direction of acceleration is 30.1 degree north of west